# Stimulus Modality Matters: Impact of Perceptual Evaluations from Different Modalities on Speech Emotion Recognition System Performance

### v2

@@ Abstract
Index Terms:
speech emotion recognition, the effects of stimulus modality, the ambiguity of emotions
@@ III-B Proposed All-inclusive Label Set and Rationales
We propose an innovative approach for training SER systems. This approach involves creating an all-inclusive label set (All) that integrates labels elicited from uni-modal and multi-modal emotional stimuli. By leveraging the comprehensive spectrum of emotional cues available from audio, visual, and audio-visual sources, we aim to harness the synergistic effect of multi-modal input to improve the performance of SER systems. This strategy is directly inspired by the inherent human capability to more accurately perceive and interpret emotions when multiple modal cues are available.
In the example of Table I, we summarize how to convert the raw annotations into the training/testing labels generated by the different modalities. The all-inclusive label (All) considers all labels elicited by voice-only, facial-only, and audio-visual stimuli. We consider a six-class emotion task, including anger (A), disgust (D), fear (F), happiness (H), sadness (S), and a neutral state (N). For voice-only elicitation, the annotations are {A*2, S*2, N*4}; for facial-only elicitation, the annotations are {A*4, S*2, N*2}; and for audio-visual elicitation, the annotations are {A*6, S*2}. The proposed all-inclusive label integrates all ratings, resulting in annotations of {A*12, S*6, N*6}. It is important to note that the labels are distributional labels for the training stage, and they are converted into binary vectors when the values are higher than the defined threshold described in Section IV-C, which is 1/C, where C is the number of emotions. In this work, we have six emotions in total (threshold = 1/6), so the testing label of the audio-visual scenario (1,1,0,0,0,0) is different from others (1,1,0,0,0,1) after applying the threshold method introduced in [10]. We follow [17] to allow the samples to have more than one emotion to reflect the nature of emotion perception that could involve mixed emotions from the psychology perspective [18].
@@ V Results and Analysis
TABLE IV: Overview of SER performances based on the WavLM Large in the cross-testing conditions. The AV represents “Audio-Visual”. For column, Overall, we also indicate the lower and upper bound of the confidence interval between 2.75% and 97.5% for each result (lower bound, upper bound) using [40] in macro-F1 scores. For other columns, we use sample-based
F1-scores. Train Set Test Set Overall (lower bound, upper bound) Angry Sad Disgust fear Neutral Happy Voice Voice 0.7117 (0.7043,0.7185) 0.7738 0.6285 0.6301 0.6665 0.9206 0.6508 Face 0.6146 (0.6079,0.6215) 0.6979 0.5583 0.5804 0.5586 0.7488 0.5436 AV 0.6486 (0.6417,0.6548) 0.7394 0.6154 0.6115 0.6289 0.7173 0.5792 All 0.6873 (0.6803,0.6938) 0.7419 0.6428 0.6394 0.6242 0.8694 0.6059 Voice Face 0.5634 (0.5555,0.5703) 0.6139 0.5075 0.5233 0.4804 0.6713 0.5841 Face 0.6366 (0.6303,0.6433) 0.6597 0.5458 0.6061 0.5835 0.7414 0.6834 AV 0.6382 (0.6320,0.6445) 0.6559 0.5337 0.6052 0.5775 0.7415 0.7154 All 0.6406 (0.6340,0.6471) 0.6548 0.5572 0.6152 0.5662 0.7201 0.7301 Voice AV 0.6418 (0.6350,0.6483) 0.7164 0.6155 0.6100 0.5802 0.6949 0.6337 Face 0.6750 (0.6691,0.6812) 0.7484 0.5593 0.6507 0.6264 0.7628 0.7021 AV 0.7076 (0.7016,0.7138) 0.7650 0.6246 0.6705 0.6679 0.7742 0.7436 All 0.7085 (0.7025,0.7142) 0.7539 0.6398 0.6936 0.6442 0.7534 0.7662
@@
Which labels elicited by various modalities are the most effective for SER systems? We choose the best model (WavLM Large) in Table III as the backbone model for the experiments. Table IV summarizes the performance of SER systems trained with labels elicited by various modalities and evaluated on test sets defined by labels elicited by different modalities. The “Train Set” column shows the models trained by which label type (Voice, Face, AV, or All), and the “Test Set” column shows the testing label type that decides the ground truth of the test sets. The testing conditions are of three types: Voice, Face, and AV. In the Overall column of Table IV, we report the macro-F1 scores along with the lower and upper bounds of the confidence interval between 2.75% and 97.5% for each result. Additionally, we present the F1 scores for the recognition performance of each emotion.

### v4

@@ III-B Proposed All-inclusive Label Set and Rationales
Fig. 1 illustrates our proposed innovative approach for training SER systems, which involves creating an all-inclusive label set (All) that integrates labels derived from uni-modal and multi-modal emotional stimuli. By leveraging the comprehensive spectrum of emotional cues available from audio, visual, and audio-visual sources, we aim to harness the synergistic effect of multi-modal input to improve the performance of SER systems. This strategy is directly inspired by the inherent human capability to more accurately perceive and interpret emotions when multiple modal cues are available.
In the example of Table I, we summarize how to convert the raw annotations into the training/testing labels generated by the different modalities. The all-inclusive label (All) considers all labels elicited by voice-only, facial-only, and audio-visual stimuli. We consider a six-class emotion task, including anger (A), disgust (D), fear (F), happiness (H), sadness (S), and a neutral state (N). For voice-only elicitation, the annotations are {A*2, S*2, N*4}; for facial-only elicitation, the annotations are {A*4, S*2, N*2}; and for audio-visual elicitation, the annotations are {A*6, S*2}. The proposed all-inclusive label integrates all ratings, resulting in annotations of {A*12, S*6, N*6}.
Importantly, the labels used during the training stage are distributional and are converted into binary vectors when their values surpass the defined threshold outlined in Section IV-C, which is 1/C, where C is the number of emotions. In this work, we have six emotions in total (threshold = 1/6), so the testing label of the audio-visual scenario (1,1,0,0,0,0) is different from others (1,1,0,0,0,1) after applying the threshold method introduced in [10]. We follow [17] to allow the samples to have more than one emotion to reflect the nature of emotion perception that could involve mixed emotions from the psychology perspective [18].
@@ V Results and Analysis
TABLE IV: Overview of SER performances based on the WavLM Large in the cross-testing conditions. The AV represents “Audio-Visual”. For column, Overall, we also indicate the lower and upper bound of the confidence interval between 2.75% and 97.5% for each result (lower bound, upper bound) using [40] in macro-F1 scores. For other columns, we use sample-based F1-scores.
Train Set Test Set Overall (lower bound, upper bound) Angry Sad Disgust Fear Neutral Happy
Voice Voice 0.7117 (0.7043,0.7185) 0.7738 0.6285 0.6301 0.6665 0.9206 0.6508
Face 0.6146 (0.6079,0.6215) 0.6979 0.5583 0.5804 0.5586 0.7488 0.5436
AV 0.6486 (0.6417,0.6548) 0.7394 0.6154 0.6115 0.6289 0.7173 0.5792
All 0.6873 (0.6803,0.6938) 0.7419 0.6428 0.6394 0.6242 0.8694 0.6059
Voice Face 0.5634 (0.5555,0.5703) 0.6139 0.5075 0.5233 0.4804 0.6713 0.5841
Face 0.6366 (0.6303,0.6433) 0.6597 0.5458 0.6061 0.5835 0.7414 0.6834
AV 0.6382 (0.6320,0.6445) 0.6559 0.5337 0.6052 0.5775 0.7415 0.7154
All 0.6406 (0.6340,0.6471) 0.6548 0.5572 0.6152 0.5662 0.7201 0.7301
Voice AV 0.6418 (0.6350,0.6483) 0.7164 0.6155 0.6100 0.5802 0.6949 0.6337
Face 0.6750 (0.6691,0.6812) 0.7484 0.5593 0.6507 0.6264 0.7628 0.7021
AV 0.7076 (0.7016,0.7138) 0.7650 0.6246 0.6705 0.6679 0.7742 0.7436
All 0.7085 (0.7025,0.7142) 0.7539 0.6398 0.6936 0.6442 0.7534 0.7662
@@
Which labels elicited by various modalities are the most effective for SER systems? We choose the best model (WavLM Large) in Table III as the backbone model for the experiments. Table IV summarizes the performance of SER systems trained with labels elicited by various modalities and evaluated on test sets defined by labels elicited by different modalities. The “Train Set” column shows the models trained by which label type (Voice, Face, AV, or All), and the “Test Set” column shows the testing label type that decides the ground truth of the test sets. The testing conditions are of three types: Voice, Face, and AV. In the Overall column of Table IV, we report the macro-F1 scores along with the lower and upper bounds of the confidence interval between 2.75% and 97.5% for each result. Additionally, we present the sample-based F1 scores for the recognition performance of each emotion.